Derek plans to retire on his 65th birthday. However, he plans to work part-time until he turns 71.00. During these years of part-time work, he will neither make deposits to nor take withdrawals from his retirement account. Exactly one year after he fully retires on the day he turns 71.0, he will begin to make annual withdrawals of $170,135.00 from his retirement account until he turns 85.00. He he will make contributions to his retirement account from his 26th birthday to his 65th birthday. To reach his goal, what must the contributions be? Assume a 4.00% interest rate. SHOW FINANCE CALCULATOR KEY STROKES
Answer #1

Assume that Derek is now 25 years old.

Step 1: Find the amount needed at age 71

Using financial calculator

Input:

PMT= 170135

N = 85-71 = 14

I/Y = 4

Hence find PV as 1,797,156.92

Step 2: Find PV of the amount at age 65

FV=1,797,156.92

I/Y=4

N= 71-65 = 6

Hence find PV as 1,420,319.22

Step 3: Find contributions

FV= 1,420,319.22

I/Y= 4

N= 65-25=40

Hence findPMT as 14,946.71

Annual contributions to be made are $14,946.71